which number completes the square of the expression 2x^ + 3x = 0?
A.9
B.9/8
C.9/16
D.-9/2

Answer by stanbon(57984)

which number completes the square of the expression
2x^2 + 3x = 0?
---
2(x^2+(3/2)x+(3/4)^2 = 2(3/4)^2
-------
Ans: (3/4)^2 = 9/16
========================
Cheers,
Stan H.
===========
A.9
B.9/8
C.9/16
D.-9/2